The length of a rectangle is 4 cubits more than twice its width. Find the dimensions of the rectangle if the area of the rectangle is 126 square cubits.
Answer by flame8855(424)

let l be the length
let w be the width
l= 4 +2w
lw=126
w(4+2w)=126
4w + 2w^2 = 126
w = 7
l = 18